OpenStudy (anonymous):

g(x) = integral from 3 to x e^((5t)^2-2t)) find the derivative i dont know how to do this at all if anyone can help

OpenStudy (tkhunny):

If \(g(x) = \int\limits_{a}^{b}f(t)\;dt = F(b) - F(a)\) if the antiderivative exists. Then \(g'(x) = F'(b) - F'(a) = f(b) - f(a)\)